OpenStudy (anonymous):

Find the value of y when x equals 1 in this equation. 2x + (-7y) = -54

OpenStudy (anonymous):

2x + (-7y) = -54 2(1) -7y = -54 2 -7y = -54 -7y = -56 y = -56/-7 y = 8

OpenStudy (anonymous):

8

OpenStudy (anonymous):

y=8